# Krylov complexity as an order parameter for quantum chaotic-integrable transitions

### v2

@@ Abstract
Krylov complexity has recently emerged as a new paradigm to characterize quantum chaos in many-body systems. However, which features of Krylov complexity are prerogative of quantum chaotic systems and how they relate to more standard probes, such as spectral statistics or out-of-time-order correlators (OTOCs), remain open questions. Recent insights have revealed that in quantum chaotic systems Krylov state complexity exhibits a distinct peak during time evolution before settling into a well-understood late-time plateau. In this work, we propose that this Krylov complexity peak (KCP) is a hallmark of quantum chaotic systems and suggest that its height could serve as an ‘order parameter’ for quantum chaos. We demonstrate that the KCP effectively identifies chaotic-integrable transitions in the mass-deformed Sachdev-Ye-Kitaev model at both infinite and finite temperature, aligning with results from spectral statistics and OTOCs. Our findings offer a new diagnostic tool for quantum chaos that is operator-independent, potentially leading to more ‘universal’ insights and a deeper understanding of general properties in quantum chaotic systems.
@@
Introduction. Chaos is a widespread phenomenon in nature. While substantial progress has been made in understanding classical chaos Hilborn (2000), its definition and characterization in the quantum realm, particularly in many-body systems, remain significantly less understood.
@@
Traditionally, quantum chaos has been linked to the Bohigas-Giannoni-Schmit (BGS) conjecture Bohigas et al. (1984); Bohigas (1984); Guhr et al. (1998); Bohigas and Giannoni (2005), asserting that the energy spectra of quantum systems with chaotic classical counterparts match the statistical predictions of random matrix theory (RMT). Specifically, quantum chaotic systems are expected to display RMT features such as level repulsion and spectral rigidity Bohigas et al. (1984); Berry (1985); Muller et al. (2004), which are therefore accepted as fingerprints of late-time quantum chaos in many-body systems.
@@
Conversely, in quantum chaotic systems at early times, the time evolution of specific out-of-time-order correlators (OTOCs) exhibits a phase of exponential growth Larkin and Ovchinnikov (1969); Berman and Zaslavsky (1978), governed by a non-zero Lyapunov exponent $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena et al. (2016). This behavior serves as an additional indicator of quantum chaos at complementary time scales.
@@
Modern explorations of quantum chaos have prominently featured both level statistics and OTOCs, bolstered by intriguing links between many-body chaos and quantum gravitational systems Shenker and Stanford (2014, 2015); García-García and Verbaarschot (2016); Cotler et al. (2017a); de Boer et al. (2018); Stanford and Witten (2020). In this context, Krylov complexity Parker et al. (2019); Balasubramanian et al. (2022) has recently emerged as a new valuable tool for characterizing quantum chaos, providing an alternative diagnostic beyond traditional analyses. It has been employed in RMT Balasubramanian et al. (2022, 2023); Tang (2023); Caputa et al. (2024); Bhattacharjee and Nandy (2024) and many other quantum chaotic systems, including quantum billiards Hashimoto et al. (2023); Camargo et al. (2024a); Balasubramanian et al. (2024), spin chains Rabinovici et al. (2022a); Scialchi et al. (2024a); Gill et al. (2024); Bhattacharya et al. (2024); Camargo et al. (2024b); Scialchi et al. (2024b), and various flavors of the Sachdev–Ye–Kitaev (SYK) model Rabinovici et al. (2021); Bhattacharjee et al. (2023a); Hörnedal et al. (2022); Erdmenger et al. (2023a); Chapman et al. (2024). Additionally, Krylov complexity has been discussed in several other contexts including topological and quantum phase transitions Caputa and Liu (2022); Afrasiar et al. (2022); Caputa et al. (2023); Pal et al. (2023), quantum batteries Kim et al. (2022), bosonic systems describing ultra-cold atoms Bhattacharyya et al. (2023), saddle-dominated scrambling Bhattacharjee et al. (2022); Huh et al. (2024), and open quantum systems Bhattacharya et al. (2022); Bhattacharjee et al. (2023b); Mohan (2023); Bhattacharya et al. (2023); Bhattacharjee et al. (2024); Carolan et al. (2024) among others – see Nandy et al. (2024) for a comprehensive review. Two forms of Krylov complexity have been proposed in the literature: the original type, which addresses operator growth Parker et al. (2019), and a newer version, which evaluates the spread of a time-evolving quantum state within a specific subspace of the Hilbert space Balasubramanian et al. (2022). The latter will be the focus of this Letter.
@@
For time-evolved thermofield double (TFD) states in RMT (see Camargo et al. (2024b) for a discussion on state-dependence of Krylov complexity), Krylov state complexity exhibits four distinct phases: an initial linear ramp, a peak, a subsequent decline, and a plateau. According to Balasubramanian et al. (2022), the peak overshooting the plateau followed by a decline appears to be a universal characteristic of quantum chaotic many-body systems and is therefore expected to be absent in integrable systems.
@@
Building on these observations, this Letter provides further evidence that the Krylov complexity peak (KCP) is a defining feature of quantum chaos, potentially serving as an ‘order parameter’ for quantum chaotic phases. Simply put, we propose that the KCP vanishes in integrable systems and that its height exhibits critical dynamics, capable of diagnosing quantum chaotic to integrable transitions at both infinite and finite temperature.
@@
To provide evidence for our proposal, we turn to the SYK model and its variants, which serve as useful toy models. Specifically, we focus on the mass-deformed SYK model Song et al. (2017); Eberlein et al. (2017); García-García et al. (2018), which has been extensively studied for its relevance to quantum chaotic to integrable transitions, and has been thoroughly characterized through both level statistics and properties of the OTOCs García-García et al. (2018); Nosaka et al. (2018); Kim and Cao (2021a); García-García et al. (2021); Lunkin et al. (2020); Nandy et al. (2022); Menzler and Jha (2024).
@@
Deformed SYK model. We consider the mass-deformed SYK model García-García et al. (2018), which involves $N$ fermions in $0+1$ dimensions. This model extends the original SYK model Chowdhury et al. (2022) by including an additional quadratic term, known as the random mass term, alongside the random quartic interactions. The Hamiltonian of the model is
@@
Overall, the deformed SYK model provides an ideal playground to test which features of Krylov complexity are unique to quantum chaotic systems and how they evolve as the system transitions to an integrable phase.
Krylov complexity and spectral form factor. The determination of Krylov complexity involves constructing the Krylov basis $\{|K_{n}\rangle\}$, achieved through the so-called Lanczos algorithm  Lanczos (1950); Viswanath and Müller (1994). This procedure yields two sets of Lanczos coefficients, $\{a_{n},\,b_{n}\}$, that encapsulate all information regarding the dynamics of the state. These coefficients correspond to the tridiagonal elements of the Hamiltonian when expressed in the Krylov basis:
@@
As the initial state, we consider the TFD state,
@@
where $Z(\beta)=\sum_{n}e^{-\beta E_{n}}$ is the partition function at inverse temperature $\beta$.
@@
Given the definitions provided in Eq.(4) and Eq.(6), it is natural to assume that a direct link between Krylov complexity and the SFF might exist. In fact, considering the Krylov complexity of the TFD state, the SFF can be interpreted as the survival probability of the time-evolved TFD state del Campo et al. (2017); Caputa et al. (2024): $\textrm{SFF}(t)=|\psi_{0}(t)|^{2}$. Additionally, for the TFD state with $\beta=0$, which is a maximally-entangled state, the late-time behavior of the Krylov complexity obeys the following identity Cotler et al. (2017a); Rabinovici et al. (2021, 2022b); Erdmenger et al. (2023a)
@@
where $C(t=\infty)={(d-1)}/{2}$, with $d$ being the system size. This identity provides a non-local (in time) constraint relating SFF$(t)$ to $C(t)$, reminiscent of a sum rule.
Results. We calculate the time-dependent Krylov complexity, as defined in Eq.(4), for the mass-deformed SYK model as a function of the mass parameter $\kappa$ and the inverse temperature $\beta$. We perform our computations with a system size of $N=26$ and express all dimensionful quantities in units of the coupling $J$, following Ref.García-García et al. (2018). As noted in Balasubramanian et al. (2022), $N=26$ is sufficiently large to ensure convergence of the numerical results. Therefore, finite-size effects on Krylov complexity are minimal for this choice of $N$. For further details on the numerical methods, we refer the reader to the Supplemental Material.
@@
As $\beta$ increases and we move away from the infinite temperature limit, three key phenomena are observed. (I) In the quantum chaotic phase (small $\kappa$), the value of $\Delta C$ decreases, indicating that the height of the peak in $C(t)$ relative to the late-time plateau value becomes temperature-dependent. This behavior is consistent with the results for $\kappa=0$ reported in Balasubramanian et al. (2022). (II) The point of continuous transition from $\Delta C\neq 0$ to $\Delta C=0$ shifts to lower values of $\kappa$. (III) The width of this transition increases, with the KCP order parameter exhibiting a smooth crossover rather than a sharp critical transition.
@@
Using the empirical function from Eq.(10) and taking $\kappa_{c}(0)\approx 66$, we estimate that for $\kappa=1$, the critical value of $\beta$ separating chaotic and integrable phases is approximately $\beta_{c}\approx 6.58$. This value lies in between the blue and purple lines in Fig. 3. Our numerical results show that this prediction, derived from the new KCP order parameter, matches the observed disappearance of the linear ramp in SFF$(t)$. This agreement not only verifies that the presence of a linear ramp and corresponding dip are clear indicators of quantum chaotic systems, which disappear as the system becomes integrable, but it also independently supports the validity and utility of the order parameter $\Delta C$ as a reliable benchmark for assessing quantum chaotic behavior.
@@
These observations hint at a deeper relationship between Krylov complexity $C(t)$ and the SFF$(t)$ that extends beyond the constraints described in Eq. (8). As noted in Erdmenger et al. (2023b), a potential connection between the KCP and the linear ramp observed in the SFF seems to be emerging. Additionally, variations of the SFF explored in previous studies García-García et al. (2018); Nosaka et al. (2018) merit further investigation, particularly regarding their relation and potential interplay with Krylov complexity.

### v4

@@ Abstract
Krylov complexity has recently emerged as a new paradigm to characterize quantum chaos in many-body systems. However, which features of Krylov complexity are prerogative of quantum chaotic systems and how they relate to more standard probes, such as spectral statistics or out-of-time-order correlators (OTOCs), remain open questions. Recent insights have revealed that in quantum chaotic systems Krylov state complexity exhibits a distinct peak during time evolution before settling into a well-understood late-time plateau. In this work, we propose that this Krylov complexity peak (KCP) is a hallmark of quantum chaotic systems and suggest that its height could serve as an ‘order parameter’ for quantum chaos. We demonstrate that the KCP effectively identifies chaotic-integrable transitions in two representative quantum mechanical models at both infinite and finite temperature: the mass-deformed Sachdev-Ye-Kitaev model and the sparse Sachdev-Ye-Kitaev model. Our findings align with established results from spectral statistics and OTOCs, while introducing an operator-independent diagnostic for quantum chaos, offering more ‘universal’ insights and a deeper understanding of the general properties of quantum chaotic systems.
@@ I Introduction
Chaos is a widespread phenomenon in nature. While substantial progress has been made in understanding classical chaos Hilborn (2000), its definition and characterization in the quantum realm, particularly in many-body systems, remain significantly less understood.
@@
Traditionally, quantum chaos has been linked to the Bohigas-Giannoni-Schmit (BGS) conjecture Bohigas et al. (1984); Bohigas (1984); Guhr et al. (1998); Bohigas and Giannoni (31 May 2005), asserting that the energy spectra of quantum systems with chaotic classical counterparts match the statistical predictions of random matrix theory (RMT). Specifically, quantum chaotic systems are expected to display RMT features such as level repulsion and spectral rigidity Bohigas et al. (1984); Berry (1985); Muller et al. (2004), which are therefore accepted as fingerprints of late-time quantum chaos in many-body systems.
@@
Conversely, in quantum chaotic systems with many degrees of freedom, ranging from the SYK model to black holes and other large-$N$ systems, at early times, the time evolution of specific out-of-time-order correlators (OTOCs) exhibits a phase of exponential growth Larkin and Ovchinnikov (1969); Berman and Zaslavsky (1978), governed by a non-zero Lyapunov exponent $\lambda_{L}\leq 2\pi k_{B}T/\hbar$ Maldacena et al. (2016). This behavior serves as an additional indicator of quantum chaos at complementary time scales.
@@
Modern explorations of quantum chaos have prominently featured both level statistics and OTOCs, bolstered by intriguing links between many-body chaos and quantum gravitational systems Shenker and Stanford (2014, 2015); García-García and Verbaarschot (2016); Cotler et al. (2017a); de Boer et al. (2018); Stanford and Witten (2020). In this context, Krylov complexity Parker et al. (2019); Balasubramanian et al. (2022) has recently emerged as a new valuable tool for characterizing quantum chaos, providing an alternative diagnostic beyond traditional analyses. It has been employed in RMT Balasubramanian et al. (2022, 2023); Tang (2023); Caputa et al. (2024); Bhattacharjee and Nandy (2024) and many other quantum chaotic systems, including quantum billiards Hashimoto et al. (2023); Camargo et al. (2024a); Balasubramanian et al. (2024), spin chains Rabinovici et al. (2022a); Scialchi et al. (2024a); Gill et al. (2024); Bhattacharya et al. (2024); Camargo et al. (2024b); Scialchi et al. (2024b), and various flavors of the Sachdev–Ye–Kitaev (SYK) model Rabinovici et al. (2021); Bhattacharjee et al. (2023a); Hörnedal et al. (2022); Erdmenger et al. (2023a); Chapman et al. (2024). Additionally, Krylov complexity has been discussed in several other contexts including topological and quantum phase transitions Caputa and Liu (2022); Afrasiar et al. (2022); Caputa et al. (2023); Pal et al. (2023), quantum batteries Kim et al. (2022), bosonic systems describing ultra-cold atoms Bhattacharyya et al. (2023), saddle-dominated scrambling Bhattacharjee et al. (2022); Huh et al. (2024), and open quantum systems Bhattacharya et al. (2022); Bhattacharjee et al. (2023b); Mohan (2023); Bhattacharya et al. (2023); Bhattacharjee et al. (2024); Carolan et al. (2024) among others – see Nandy et al. (2024) for a comprehensive review. Two forms of Krylov complexity have been proposed in the literature: the original type, which addresses operator growth Parker et al. (2019), and a newer version, which evaluates the spread of a time-evolving quantum state within a specific subspace of the Hilbert space Balasubramanian et al. (2022). The latter will be the focus of this manuscript.
@@
For time-evolved thermofield double (TFD) states in RMT (see Camargo et al. (2024b) for a discussion on state-dependence of Krylov complexity), Krylov state complexity exhibits four distinct phases: an initial linear ramp, a peak, a subsequent decline, and a plateau. According to Balasubramanian et al. (2022), the peak overshooting the plateau followed by a decline appears to be a universal characteristic of quantum chaotic many-body systems and is therefore expected to be absent in integrable systems. In this context, the importance of the TFD state is motivated not only by its role in its connection between Krylov complexity and other chaos probes, such as the spectral form factor (discussed in detail later in the text), but also for its potential relevance in holography, where it serves as the holographic dual of a two-sided black hole.
@@
Building on these observations, this paper provides further evidence that the Krylov complexity peak (KCP) is a defining feature of quantum chaos, potentially serving as an ‘order parameter’ for quantum chaotic phases. Simply put, we propose that the KCP vanishes in integrable systems and that its height exhibits critical dynamics, capable of diagnosing quantum chaotic to integrable transitions at both infinite and finite temperature.
@@
To provide evidence for our proposal, we turn to the SYK model and its variants, which serve as useful toy models. Specifically, we focus on two representative examples: the mass-deformed SYK Song et al. (2017); Eberlein et al. (2017); García-García et al. (2018) and the sparse SYK Xu et al. (2020); García-García et al. (2021) models, which have been extensively studied for their relevance to quantum chaotic to integrable transitions, and have been thoroughly characterized through both level statistics and properties of the OTOCs Song et al. (2017); Eberlein et al. (2017); García-García et al. (2018); Xu et al. (2020); García-García et al. (2021); Nosaka et al. (2018); Kim and Cao (2021a); García-García et al. (2021); Lunkin et al. (2020); Nandy et al. (2022); Menzler and Jha (2024); Cáceres et al. (2022); Orman et al. (2024); Cáceres et al. (2023); García-García et al. (2024).
This paper is structured as follows: Section II introduces the SYK models under consideration. Section III provides an overview of Krylov complexity and the spectral form factor. In Section IV, we show that Krylov complexity effectively identifies chaotic-integrable transitions in these models, in line with the appearance and disappearance of the ramp in the spectral form factor. Finally, Section V presents our main conclusions
II Computational Models
II.1 Mass-deformed SYK model
@@
As a first example, we consider the mass-deformed SYK model García-García et al. (2018), which involves $N$ fermions in $0+1$ dimensions. This model extends the original SYK model Chowdhury et al. (2022) by including an additional quadratic term, known as the random mass term, alongside the random quartic interactions. The Hamiltonian of the model is given by
@@
II.2 Sparse SYK model
As a second example, we consider the sparse SYK model Xu et al. (2020); García-García et al. (2021). In this case, the Hamiltonian is
$H=\frac{1}{4!}\sum^{N}_{i,j,k,l=1}\,x_{ijkl}\,J_{ijkl}\,\chi_{i}\,\chi_{j}\,\chi_{k}\,\chi_{l}\,,$ (2)
where $\chi_{i}$ are Majorana fermions satisfying $\{\chi_{i},\,\chi_{j}\}=\delta_{ij}$. The coupling constants $J_{ijkl}$ are Gaussian-distributed random variables with zero mean and standard deviation
$\sigma=\sqrt{\frac{6J^{2}}{pN^{3}}}\,.$ (3)
Furthermore, $x_{ijkl}$ equals 1 with probability $p$ and 0 with probability $1-p$. The parameter $p$ determines the number of non-zero terms in the Hamiltonian, $kN$, given by
$kN=p\,\binom{N}{4}\,,$ (4)
which controls the amount sparseness of the model. For $p=1$, the model reduces to the standard SYK model. As $p$ (or equivalently $k$) decreases, the model transitions towards an integrable regime, as evidenced by level statistics analysis Cáceres et al. (2022); Orman et al. (2024); García-García et al. (2021), which identifies a critical value of $k_{c}\approx 1$ for this transition. This transition is further reflected in the behavior of the Lyapunov exponent, indicated by a significant reduction in the exponential growth of OTOCs Cáceres et al. (2023). In a parallel analysis it was shown that the emergence of gravitational physics at low temperatures —specifically the onset of Schwarzian dynamics— requires $k$ to lie within the range of 1/4 to 4 Xu et al. (2020).
The Krylov complexity of states has been examined in the sparse SYK model, as discussed in Jha and Roy (2024), though from a different perspective. In contrast to their approach, we will use the TFD state as our initial condition, aligning with the original conjecture for Krylov complexity Balasubramanian et al. (2022). Additionally, we will extend the analysis to investigate the temperature dependence of this complexity.
Overall, these two variants of the SYK model serve as ideal playgrounds for exploring the features of Krylov complexity that are unique to quantum chaotic systems and examining how these features evolve as the system transitions to an integrable phase. For both models, we will take $J=1$ for our computations.
III Krylov complexity and
spectral form factor
The calculation of Krylov complexity involves constructing the Krylov basis $\{|K_{n}\rangle\}$, achieved through the so-called Lanczos algorithm  Lanczos (1950); Viswanath and Müller (1994). This procedure yields two sets of Lanczos coefficients, $\{a_{n},\,b_{n}\}$, which encode all information regarding the system’s dynamics. These coefficients correspond to the tridiagonal elements of the Hamiltonian when expressed in the Krylov basis:
@@ III Krylov complexity and
As initial state, we consider the TFD state,
@@
with $|n\rangle$ and $E_{n}$ indicating respectively the eigenstates and the eigenvalues of the Hamiltonian $H$. This TFD state is built from the tensor product of two copies of the original Hilbert space. Here, $Z(\beta)=\sum_{n}e^{-\beta E_{n}}$ is the partition function at inverse temperature $\beta$ [^1] In the Hessenberg form using Householder reflections, the standard initial state is chosen as $(1,0,0,\cdots)^{\text{T}}$. Therefore, a basis transformation is required to align the given initial vector with this standard state. For more details, see Balasubramanian et al. (2022). Recall that the TFD state evolves under ($H_{L}+H_{R}$)/2, where $H_{L}$ and $H_{R}$ act independently on the left and right copies of the Hamiltonian, respectively.
@@
Given the definitions provided in Eq.(7) and Eq.(9), it is natural to assume that a direct link between Krylov complexity and the SFF might exist. In fact, considering the Krylov complexity of the TFD state, the SFF can be interpreted as the survival probability of the time-evolved TFD state del Campo et al. (2017); Caputa et al. (2024): $\textrm{SFF}(t)=|\psi_{0}(t)|^{2}$, where $\psi_{n=0}(t)$ is determined by solving Eq.(6). Additionally, for the TFD state with $\beta=0$, which is a maximally-entangled state, the late-time behavior of the Krylov complexity obeys the following identity Cotler et al. (2017a); Rabinovici et al. (2021, 2022b); Erdmenger et al. (2023a)
@@
where $C(t=\infty)={(d-1)}/{2}$, with $d$ being the system size related to $N$ as $d=2^{N/2-1}$. [^2] In the case of the TFD state with finite $\beta$, i.e., when the system is no longer in a maximally entangled state, a simple relation as in Eq. (11) may no longer hold. For further details on this point, see Ref. Erdmenger et al. (2023a). This identity provides a non-local (in time) constraint relating SFF$(t)$ to $C(t)$, reminiscent of a sum rule.
IV Results
IV.1 Mass-deformed SYK model
We calculate the time-dependent Krylov complexity, as defined in Eq.(7), for the mass-deformed SYK model as a function of the mass parameter $\kappa$ and the inverse temperature $\beta$. We perform our computations with a system size of $N=26$ and express all dimensionful quantities in units of the coupling $J$, following Ref.García-García et al. (2018). As noted in Balasubramanian et al. (2022), $N=26$ is sufficiently large to ensure convergence of the numerical results. Therefore, finite-size effects on Krylov complexity are minimal for this choice of $N$. For further details on the numerical methods, we refer the reader to the Appendix A.
[…]
@@ IV.1 Mass-deformed SYK model
As $\beta$ increases and we move away from the infinite temperature limit, three key phenomena are observed. (I) In the quantum chaotic phase (small $\kappa$), the value of $\Delta C$ decreases, indicating that the height of the peak in $C(t)$ relative to the late-time plateau value becomes temperature-dependent. This behavior is consistent with the results for $\kappa=0$ reported in Balasubramanian et al. (2022). (II) The point of continuous transition from $\Delta C\neq 0$ to $\Delta C=0$ shifts to lower values of $\kappa$. (III) The width of this transition (e.g., the width is from $\kappa\approx 5$ to $\kappa\approx 30$ for $\beta=0$) increases, with the KCP order parameter exhibiting a smooth crossover rather than a sharp critical transition.
@@
Using the empirical function from Eq.(13) and taking $\kappa_{c}(0)\approx 66$, we estimate that for $\kappa=1$, the critical value of $\beta$ separating chaotic and integrable phases is approximately $\beta_{c}\approx 6.58$. This value lies in between the blue and purple lines in Fig. 3. Our numerical results show that this prediction, derived from the new KCP order parameter, matches the observed disappearance of the ramp in SFF$(t)$. This agreement not only verifies that the presence of a ramp and corresponding dip are clear indicators of quantum chaotic systems, which disappear as the system becomes integrable, but it also independently supports the validity and utility of the order parameter $\Delta C$ as a reliable benchmark for assessing quantum chaotic behavior.
@@
These observations hint at a deeper relationship between Krylov complexity $C(t)$ and the SFF$(t)$ that extends beyond the constraints described in Eq. (11). As noted in Erdmenger et al. (2023b), a potential connection between the KCP and the ramp observed in the SFF seems to be emerging. Additionally, variations of the SFF explored in previous studies García-García et al. (2018); Nosaka et al. (2018) merit further investigation, particularly regarding their relation and potential interplay with Krylov complexity.
Together with the disappearance of the ramp, there is an additional observation to be made. As shown in Fig. 1 (bottom), for a fixed value of $\kappa$ and varying $\beta$, there is a critical value at which the KCP disappears. For example, for $\kappa=1$ the KCP disappears in between $\beta=1$ (orange) and $\beta=3$ (green). This transition is also evident in Fig. 3, where the dip in the SSF vanishes at around the same value of $\beta$ (green curve). It is therefore plausible that the depth of the dip in the SFF may serve as an alternative order parameter to detect the chaotic to integrable transition, similar to the KCP proposed above. However, as analyzed in detail in Appendix A.4, this is not the case in general. In fact, we find that this feature is evident only at finite $\beta$, but it disappears in the limit of $\beta\rightarrow 0$. This suggests that the KCP is a more robust indicator of the chaos-integrable transition.
IV.2 Sparse SYK model
To further illustrate the robustness of the KCP as an order parameter for chaotic-integrable transitions, we investigate the sparse SYK model. As with our study of the mass-deformed SYK model, we use a system size of $N=26$ and examine Krylov complexity across different levels of sparsity, parameterized by $p$, or $k$ in Eq. (4).
The top panel of Fig. 4 illustrates the time evolution of normalized Krylov complexity for various sparsity values in the range $p\in[0.001,1]$ at $\beta=0$. As sparsity increases (i.e., as $p$ decreases), two key observations emerge: (i) the saturation value deviates from $C(t=\infty)={(d-1)}/{2}$, and (ii) the peak height diminishes.
The deviation in the saturation value arises because, in the sparse SYK model at sufficiently small $p$ (specifically, when $k<1$), a significant number of emergent discrete symmetries, including chiral symmetries, can lead to exact degeneracies in the spectrum García-García et al. (2021). These degeneracies modify the late-time behavior of Krylov complexity of the TFD state, resulting in a suppressed saturation value Erdmenger et al. (2023a); Camargo et al. (2024b). This behavior is consistent with previous analyses of Krylov complexity in the sparse SYK model using a different initial state Jha and Roy (2024).
In the bottom panel of Fig. 4, we plot the KCP order parameter $\Delta C$ as a function of $k$ for various values of $\beta$. We observe a transition to $\Delta C\approx 0$ with the critical value of $k_{c}\approx 1$ which is consistent with previous studies using spectral statistics methods such as the $r$-parameter statistics García-García et al. (2021). Furthermore, Fig. 5 corroborates this transition, demonstrating that the ramp in the spectral form factor also disappears around $k_{c}\approx 1$.
Our results from both the mass-deformed and sparse SYK models indicate that the KCP serves as an effective order parameter for chaotic-integrable transitions, consistent with findings from conventional spectral statistics methods. Notably, unlike the mass-deformed SYK model, the critical $k_{c}$ remains approximately constant at $k_{c}\approx 1$, regardless of the value of $\beta$. This observation may be linked to the level statistics analysis in García-García et al. (2021), which demonstrated that at $k\approx 1$, the system reaches a maximum level of sparsity, leading to the absence of level repulsion. [^3] For further details, see García-García et al. (2021), which elaborates on the observation that the distribution width exceeds the level spacing, with the distributions of the first ten eigenvalues being nearly identical. This suggests that in the regime $k<1$, the Hamiltonian is too sparse to maintain quantum chaotic features. In other words, the Hamiltonian loses its chaotic nature, rendering the influence of the probe (such as $\beta$ through the TFD state) negligible.
[…]
Figure 5: Spectral form factor, Eq. (9), as a function of time for various values of the parameter $k=575\,(\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{red}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$, $1.150\,(\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\textbf{orange}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$, $0.805\,(\color[rgb]{0.58,0.0,0.83}\definecolor[named]{pgfstrokecolor}{rgb}{0.58,0.0,0.83}\textbf{purple}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ for $N=26$. The black dashed line indicates the average linear growth characteristic of the SFF in the ramp regime. Top: $\beta=0$. Bottom: $\beta=5$.
V Discussion